OpenStudy (anonymous):

Solve the equation. Check the solution. -4 / x + 1 = -1 / x + 5

OpenStudy (anonymous):

can someone please explain what im supposed to do

OpenStudy (anonymous):

-4/(x+1) = -1/(x+5) 4/(x+1) = 1/(x+5) 4(x+5)=1(x+1)

OpenStudy (anonymous):

4x + 20 = 1x + 1???